# Category $\mathcal{O}$ for Lie superalgebras

### v1

@@ 8.2. Universally ${\bar{0}}$-bounded
Let ${\mathfrak{p}}$ be a principal parabolic in ${\mathfrak{g}}$. From Lemma 6.1.1, if $L(\lambda)$ is a simple module in ${\mathcal{O}}^{\mathfrak{p}}$ then this module is in ${\mathcal{O}}^{\mathfrak{p}_{{\bar{0}}}}$. Therefore, in the Grothendieck group $[L(\lambda)]=\sum_{\mu\in\Lambda_{{\bar{0}}}}n_{\lambda,\mu}[L_{{\bar{0}}}(\mu)]$. We say that the the simple modules in ${\mathcal{O}}^{\mathfrak{p}}$ are universally ${\bar{0}}$-bounded if and only if there exists a $C>0$ such that $n_{\lambda,\mu}\leq C$ for all $\lambda\in\Lambda$ and $\mu\in\Lambda_{{\bar{0}}}$.
@@
The module $\Lambda^{\bullet}({\mathfrak{g}}_{{\bar{1}}})\otimes L_{{\mathfrak{p}}_{{\bar{0}}}}(\sigma)$ is a finite-dimensional module over $U({\mathfrak{l}}_{{\bar{0}}})$. One can show that the composition multiplicities are bounded by the number of weights of $\gamma=\sigma+\mu$ where $\mu$ is a weight of $\Lambda^{\bullet}({\mathfrak{g}}_{{\bar{1}}})$, and $\gamma$ is dominant with respect to ${\mathfrak{l}}_{{\bar{0}}}$. This number is clearly bounded by the dimension of $\Lambda^{\bullet}({\mathfrak{g}}_{{\bar{1}}})$. Now for each $\gamma$, the composition multiplicities of [$Z_{{\bar{0}}}(\gamma):L_{{\bar{0}}}(\tau)]$ are bounded by Proposition 3.3.1(c). Hence, the composition factor multiplicities of the $U({\mathfrak{g}}_{{\bar{0}}})$-module in (8.2.2) is universally bounded for all $\sigma$, and thus the simple modules in ${\mathcal{O}}^{\mathfrak{p}}$ are universally ${\bar{0}}$-bounded.
@@ 8.3. Proof of Theorem D
In the Grothendieck group $[M]=\sum_{\sigma\in\Gamma}n_{\sigma}[L_{{\bar{0}}}(\sigma)]$. Since $\Gamma$ is finite we can bound $n_{\sigma}$ where $\sigma\in\Gamma$. Therefore,
@@ Appendix A Almost Simple Lie Superalgebras
In the appendix, we compare common setups in the study of finite dimensional Lie superalgebras over $\mathbb{C}$.
@@ A.1. Quasi-reductive Lie Superalgebras
The complete list (see Table A.1.1 below) of simple Lie superalgebras consists of the Cartan series and the classical simple Lie superalgebras. A finite-dimensional (not necessarily simple) Lie superalgebra $\mathfrak{g}$ is quasi-reductive (also called classical) if $\mathfrak{g}_{\bar{0}}$ is reductive and $\mathfrak{g}_{\bar{1}}$ is semisimple as a $\mathfrak{g}_{\bar{0}}$-module. A classical simple Lie superalgebra $\mathfrak{g}$ is basic if it has a nondegenerate invariant supersymmetric even bilinear form; while it belongs to the strange series if it is not basic. A basic classical Lie superalgebra $\mathfrak{g}$ is said to be of Type I if $\mathfrak{g}_{\bar{0}}=\mathfrak{g}_{0}$ and $\mathfrak{g}_{\bar{1}}=\mathfrak{g}_{1}\oplus\mathfrak{g}_{-1}$; otherwise, $\mathfrak{g}$ is said to be of Type II.
@@
Classical simple Lie superalgebras Cartan series $\begin{array}[]{l}W(n)\\ S(n),\overline{S}(n)\\ H(n)\end{array}$
@@ A.3.4. Type BC
It follows that $(\epsilon_{j}-\epsilon_{j+1})(t)=1$ and $(\delta_{j}-\delta_{j+1})(t)=1$ for $j=1,2,\dots,n-1$. Moreover, $\epsilon_{n}(t)=1$ and $22\delta_{n}(t)=$ 2, thus $\alpha(t)>0$ for all $\alpha\in\Phi_{{\bar{0}}}^{+}$.
@@ A.3.5. Type D
$t=\langle(\text{diag}\{a^{n},a^{n-1},\dots,a^{1},a^{-n},a^{-(n-1)},\dots,a^{-1}\},\text{diag}\{a^{n},a^{n-1},\dots,a^{1},a^{-n},\dots,a^{-1}\})$ (A.3.7)

### v2

@@ 8.2. Universally $\bar{0}$}-bounded
Let $\mathfrak{p}$} be a principal parabolic in $\mathfrak{g}$}. From Lemma 6.1.1, if $L(\lambda)$ is a simple module in ${\mathcal{O}}^{\mathfrak{p}}$ then this module is in ${\mathcal{O}}^{\mathfrak{p}_{{\bar{0}}}}$. Therefore, in the Grothendieck group $L(\lambda)]=\sum_{\mu\in\Lambda_{{\bar{0}}}}n_{\lambda,\mu}[L_{{\bar{0}}}(\mu)]$. We say that the the simple modules in ${\mathcal{O}}^{\mathfrak{p}}$ are universally $\bar{0}$}-bounded if and only if there exists a $C>0$ such that $n_{\lambda,\mu}\leq C$ for all $\lambda\in\Lambda$ and $\mu\in\Lambda_{{\bar{0}}}$.
@@
The module $\Lambda^{\bullet}({\mathfrak{g}}_{{\bar{1}}})\otimes L_{{\mathfrak{p}}_{{\bar{0}}}}(\sigma)$ is a finite-dimensional module over $U({\mathfrak{l}}_{{\bar{0}}})$. One can show that the composition multiplicities are bounded by the number of weights of $\gamma=\sigma+\mu$ where $\mu$ is a weight of $\Lambda^{\bullet}({\mathfrak{g}}_{{\bar{1}}})$, and $\gamma$ is dominant with respect to ${\mathfrak{l}}_{{\bar{0}}}$. This number is clearly bounded by the dimension of $\Lambda^{\bullet}({\mathfrak{g}}_{{\bar{1}}})$. Now for each $\gamma$, the composition multiplicities of $Z_{{\bar{0}}}(\gamma):L_{{\bar{0}}}(\tau)]$ are bounded by Proposition 3.3.1(c). Hence, the composition factor multiplicities of the $U({\mathfrak{g}}_{{\bar{0}}})$-module in (8.2.2) is universally bounded for all $\sigma$, and thus the simple modules in ${\mathcal{O}}^{\mathfrak{p}}$ are universally $\bar{0}$}-bounded.
@@ 8.3. Proof of Theorem D
In the Grothendieck group $M]=\sum_{\sigma\in\Gamma}n_{\sigma}[L_{{\bar{0}}}(\sigma)]$. Since $\Gamma$ is finite we can bound $n_{\sigma}$ where $\sigma\in\Gamma$. Therefore,
@@ Appendix A Almost Simple Lie Superalgebras
In the appendix, we compare common setups in the study of finite-dimensional Lie superalgebras over $\mathbb{C}$.
@@ A.1. Quasi-reductive Lie Superalgebras
The complete list (see Table A.1.1 below) of simple Lie superalgebras consists of the Cartan series and the classical simple Lie superalgebras. A finite-dimensional (not necessarily simple) Lie superalgebra $\mathfrak{g}$ is quasi-reductive (also called classical) if $\mathfrak{g}_{\bar{0}}$ is reductive and $\mathfrak{g}_{\bar{1}}$ is semisimple as a $\mathfrak{g}_{\bar{0}}$-module. A classical simple Lie superalgebra $\mathfrak{g}$ is basic if it has a nondegenerate invariant supersymmetric even bilinear form; while it belongs to the strange series if it is not basic. A basic classical Lie superalgebra $\mathfrak{g}$ is said to be of Type I if $\mathfrak{g}$ is $\mathbb{Z}$}-graded with $\mathfrak{g}_{\bar{0}}=\mathfrak{g}_{0}$ and $\mathfrak{g}_{\bar{1}}=\mathfrak{g}_{1}\oplus\mathfrak{g}_{-1}$; otherwise, $\mathfrak{g}$ is said to be of Type II.
@@
Classical simple Lie superalgebras
Cartan series
$\begin{array}[]{l}W(n)\\ S(n),\overline{S}(n)\\ H(n)\end{array}$
@@ A.3.4. Type BC
It follows that $(\epsilon_{j}-\epsilon_{j+1})(t)=1$ and $(\delta_{j}-\delta_{j+1})(t)=1$ for $j=1,2,\dots,n-1$. Moreover, $\epsilon_{n}(t)=1$ and $2\delta_{n}(t)=2$, thus $\alpha(t)>0$ for all $\alpha\in\Phi_{{\bar{0}}}^{+}$.
@@ A.3.5. Type D
$t=(\text{diag}\{a^{n},a^{n-1},\dots,a^{1},a^{-n},a^{-(n-1)},\dots,a^{-1}\},\text{diag}\{a^{n},a^{n-1},\dots,a^{1},a^{-n},\dots,a^{-1}\})$ (A.3.7)